40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 29 15:47:58 2019, gautam, Update, AUX, AUX laser fiber moved from AS table to PSL table 
    Reply  Sun Mar 31 18:39:45 2019, Anjali, Update, AUX, AUX laser fiber moved from AS table to PSL table Schematic_of_experimental_setup_for_frequency_stabilisation_of_1_micron_source.png1_micron_setup_for_frequency_noise_measurement.JPG1_micron_setup_for_frequency_noise_measurement_power_levels.png
Message ID: 14504     Entry time: Sun Mar 31 18:39:45 2019     In reply to: 14501
Author: Anjali 
Type: Update 
Category: AUX 
Subject: AUX laser fiber moved from AS table to PSL table 
  • Attachment #1 shows the schematic of the experimental setup for the frequency noise measurement of 1 um laser source.

  • AUX laser will be used as the seed source and it is already coupled to a 60 m fiber (PM980). The other end of the fiber was at the AS table and we have now removed it and placed in the PSL table.

  • Attachment # 2 shows the photograph of the experimental setup. The orange line shows the beam that is coupled to the delayed arm of MZI and the red dotted line shows the undelayed path.

  • As mentioned, AUX is already coupled to the 60 m fiber and the other end of the fiber is now moved to the PSL table. This end needs to be collimated. We are planning to take the same collimator from AS table where it was coupled into before. The position where the collimator to be installed is shown in attachment #2. Also, we need to rotate the mirror (as indicated in attachment #2) to get the delayed beam along with the undelayed beam and then to combine them. As indicated in attachment #2, we can install one more photo diode to perform  balanced detection.

  • We need to decide on which photodetector to be used. It could be NF1801 or PDA255.

  • We also performed the power measurement at different locations in the beam path. The different locations at which power measurement is done is shown attachment #3

  • There is an AOM in the beam path that coupled to the delayed arm of MZI. The output beam after AOM was coupled to the zero-order port during this measurement. That is the input voltage to the AOM was at 0 V, which essentially says that the beam after the AOM is not deflected and it is coupled to the zero-order port. The power levels measured at different locations in this condition are as follows. A)282 mW B)276 mW C)274 mW D)274 mW E)273 mW F)278 mW G)278 mW H)261 mW I)263 mW J)260 mW K)131 mW L)128 mW M)127 mW N)130 mW

  • It can be seen that the power is halved from J to K. This because of a neutral density filter in the path of the beam

  • In this case, we measured a power of 55 mW at the output of the delayed fiber. We then adjusted the input voltage to the AOM driver to 1 V such that the output of AOM is coupled to the first order port. This reduced the power level in the zero-order port of AOM that is coupled to the delayed arm of the MZI. In this case we measured a power of 0.8 mW at the output of delayed fiber.

  •  We must be careful about the power level that is reaching the photodetector such that it should not exceed the damage threshold of the detector.

  • The power measured at the output of undelayed path is 0.8 mW.

  • We also must place the QWP and HWP in the beam path to align the polarisation.

Quote:

[anjali, gautam]

To facilitate the 1um MZ frequency stabilization project, I decided that the AUX laser was a better candidate than any of the other 3 active NPROs in the lab as (i) it is already coupled into a ~60m long fiber, (ii) the PSL table has the most room available to set up the readout optics for the delayed/non-delayed beams and (iii) this way I can keep working on the IR ALS system in parallel. So we moved the end of the fiber from the AS table to the SE corner of the PSL table. None of the optics mode-matching the AUX beam to the interferometer were touched, and we do not anticipate disturbing the input coupling into the fiber either, so it should be possible to recover the AUX beam injection into the IFO relatively easily.

Anjali is going to post detailed photos, beam layout, and her proposed layout/MM solutions later today. The plan is to use free space components for everything except the fiber delay line, as we have these available readily. It is not necessarily the most low-noise option, but for a first pass, maybe this is sufficient and we can start building up a noise budget and identify possible improvements.

The AUX laser remians in STANDBY mode for now. HEPA was turned up while working at the PSL table, and remains on high while Anjali works on the layout.

 

Attachment 1: Schematic_of_experimental_setup_for_frequency_stabilisation_of_1_micron_source.png  68 kB  Uploaded Sun Mar 31 19:40:35 2019  | Hide | Hide all
Schematic_of_experimental_setup_for_frequency_stabilisation_of_1_micron_source.png
Attachment 2: 1_micron_setup_for_frequency_noise_measurement.JPG  6.800 MB  Uploaded Sun Mar 31 19:41:46 2019  | Hide | Hide all
1_micron_setup_for_frequency_noise_measurement.JPG
Attachment 3: 1_micron_setup_for_frequency_noise_measurement_power_levels.png  4.535 MB  Uploaded Sun Mar 31 19:42:37 2019  | Hide | Hide all
1_micron_setup_for_frequency_noise_measurement_power_levels.png
ELOG V3.1.3-